OpenStudy (xobriley):

9. What is the value of x/y when x = -16 and y = 4 a) 4 *b) -12 c) -4 d) 12

jimthompson5910 (jim_thompson5910):

replace every x with -16 replace every y with 4 so `x/y = (-16)/(4) = ???`

rebeccaxhawaii (rebeccaxhawaii):

okay so replace -16/4

rebeccaxhawaii (rebeccaxhawaii):

or \[-16\div4\]

OpenStudy (xobriley):

so d???

jimthompson5910 (jim_thompson5910):

what is 16 over 4 equal to?

jimthompson5910 (jim_thompson5910):

whatever that result is, make it negative since we have negative over positive = negative

OpenStudy (xobriley):

-4

jimthompson5910 (jim_thompson5910):

correct

OpenStudy (xobriley):

Can you help me with one more

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (xobriley):

10.Simplify the expression. 5by -2y a) -5b/2 b) 5b/2 c) -10 by *d)-5y/2

jimthompson5910 (jim_thompson5910):

5by over -2y ??

jimthompson5910 (jim_thompson5910):

is `5by over -2y` the original expression ?

OpenStudy (xobriley):

original

jimthompson5910 (jim_thompson5910):

notice how the y terms will cancel

jimthompson5910 (jim_thompson5910):

and how positive over negative = negative

jimthompson5910 (jim_thompson5910):

so you'll end up with `-5b/2`

OpenStudy (xobriley):

1. What is the simplified form of the expression? 3t2 + 8m + 4t2 + 5m - 2m a) 3t2 + 13m *b) 7t2 + 11m c) 7t2 + 15m d) 4t2 + 11m 2. Simplify the following expression. *5(9x - 4) a) 14x b) 14x - 9 c) 45x - 20 d) 45x + 20 3.Which property is illustrated by the statement? 3+5x = 5x+3 *a) commutative property of addition b) associative property c) distributive property 4. To which subset(s) of the real numbers does the number -76 belong? a) rational b) irrational *c) integers, whole numbers, rational numbers d) integers, rational numbers 5. What is the algebraic expression for for the following word phrase: a) r+8 b) r/8 *c) 8r Name the quotient of r and 8? d) r-8 6. What is the difference? -9-7 a) -2 b) 2 *c)-16 d) 16 7. What is the product? 7 (-3) a) -10 b) 4 c) 21 *d) -21 8. Simplify the expression. (3/5)4 a) 1 b) 12/20 *c) 81/625 d) 7/9 9. What is the value of x/y when x = -16 and y = 4 a) 4 b) -12 *c) -4 d) 12 10.Simplify the expression. 5by -2y a) -5b/2 b) 5b/2 c) -10 by d)-5y/2

OpenStudy (xobriley):

Got all them right?

jimthompson5910 (jim_thompson5910):

what did you get for #2

OpenStudy (xobriley):

the first one

jimthompson5910 (jim_thompson5910):

#2 Distribute 5(9x - 4) = 5*(9x) + 5(-4) = 45x - 20

jimthompson5910 (jim_thompson5910):

#4 is wrong because the set of whole numbers is this set W = {0,1,2,3,4,5,6,....} basically all the positive whole numbers and 0

OpenStudy (xobriley):

So whats the answer

jimthompson5910 (jim_thompson5910):

if you want to talk about negative numbers without decimal parts, you refer to the set of integers Z = {..., -5, -4, -3, -2, -1, 0, 1, 2, 3, 4, 5, ...}

OpenStudy (xobriley):

Rational?

jimthompson5910 (jim_thompson5910):

-76 is rational and an integer

jimthompson5910 (jim_thompson5910):

#5 is wrong too

OpenStudy (xobriley):

whats that one?

jimthompson5910 (jim_thompson5910):

quotient means result after dividing

OpenStudy (xobriley):

r\8

jimthompson5910 (jim_thompson5910):

r/8

jimthompson5910 (jim_thompson5910):

everything else looks good

OpenStudy (xobriley):

Thank you so much

jimthompson5910 (jim_thompson5910):

no problem

OpenStudy (xobriley):

What is the correct Order of Operations?

OpenStudy (xobriley):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

PEMDAS http://www.webquest.hawaii.edu/kahihi/mathdictionary/images/PEMDAS.png